# AgriFM: A Multi-source Temporal Remote Sensing Foundation Model for Crop Mapping

### v2

@@ 2.4 Paddy Rice Mapping in Monsoon Asia
2.5 Winter Wheat Mapping in Asia
@@
To comprehensively evaluate AgriFM’s performance, we conduct comparisons across three distinct methodological categories. First, we examine conventional deep learning architectures widely adopted in crop mapping: (1) A CNN (VGG16 variant (Simonyan and Zisserman, 2014)) treating temporal observations as additional input channels; (2) A CNN-LSTM hybrid extracting spatial features per timestep using VGG16 followed by temporal modeling through LSTM; and (3) A 3D CNN processing spatiotemporal cubes (height × width × time × bands). All baselines employ identical decoder structures and training protocols to isolate architectural differences, ensuring fair comparison of their spatiotemporal representation capabilities.
Second, we compare against ViT-based foundation models that incorporate temporal information during pretraining. Specifically, we evaluate SatMAE (Cong et al., 2022) (using Landsat-8 data) and Prithvi (Jakubik et al., 2023) (trained on HLS30 data), which represent the most temporally-aware open-source RSFMs currently available, despite their limited sequence lengths. These comparisons reveal how feature fusion strategies and hierarchical representations impact mapping precision, particularly for crops with subtle spectral-temporal signatures.
Third, we analyze PIS, a Swin Transformer-based foundation model lacking explicit temporal processing. While sharing similar backbone architecture and pretraining data scales with our approach, PIS’s non-temporal design helps isolate the benefits of our temporal-aware pretraining strategy. Although PIS uses the Swin transformer as its backbone, we can reuse some weights as pre-trained weights of the video Swin transformer and adopt the same architecture as our method for comparison.
2.6 Implementation details
@@
For downstream crop mapping tasks, we employ the same network structure, that is, the unified crop mapping framework that we construct. The learning rate is set to 6e-5. All experiments are conducted on four L40 GPUs, with a batch size of 16, and last for 50 epochs. We select the model that performs best on the validation dataset as our final result.
3 Cropland Mapping and Field Boundary Delineation
@@
As detailed in Tables 3 and 4, we systematically evaluate three methodological categories: (1) CNN variants (CNN, CNN-LSTM, 3DCNN) representing conventional deep learning approaches; (2) ViT-based methods (SatMAE, Prithvi) as current mainstream remote sensing foundation models pretrained with temporal data; and (3) Swin transformer (PIS, AgriFM-scratch) demonstrating hierarchical feature extraction’s impact on crop mapping. Notably, PIS’s backbone is Swin transformer and doesn’t adopt temporal data. We adaptively rearrange its partial parameters to initialize our Video Swin Transformer, enabling comparison of temporal-agnostic pretraining effects. AgriFM-scratch serves as the randomly initialized control. The ”Positive” column specifies cropland / boundary-specific metrics, while ”Average” incorporates background class performance. Figure 5 visually contrasts core metrics (Precision / Recall / F1) across methods, and Figure 7 showcases actual mapping outputs where green denotes cropland and red indicates boundaries.
Table 3: Performance comparison on cropland mapping task (%). Positive class metrics focus on cropland identification.
Model Positive (Cropland) Average
@@
AgriFM 79.35 84.63 81.90 82.68 82.87 82.70 82.74
Table 4: Performance comparison on field boundary delineation task (%). Metrics computed at field boundary pixels.
Model Positive (Boundary) Average
@@
AgriFM 73.93 73.63 73.78 79.02 78.97 78.99 80.29
For cropland mapping (Table 3), CNN architectures demonstrate strong baseline performance, with CNN-LSTM leading at 80.39% F1-score due to its explicit spatiotemporal fusion design. 3DCNN achieves only 77.93% F1 despite 3D convolutions. This shows that simply adding 3D convolution does not guarantee better spatiotemporal modeling, and also shows that the 3D convolution in traditional deep learning methods for processing spatiotemporal data may have insufficient information extraction capabilities for crop mapping tasks. ViT-based SatMAE and Prithvi trail significantly at 76.80% and 73.69% F1 respectively, confirming ViT’s architectural constraints—the 16×16 patch embedding causes irreversible spatial detail loss (evident in Figure 7’s blurred boundaries). While increasing input resolution could theoretically mitigate this issue, the computational cost becomes prohibitive for temporal sequences while distorting original resolution relationships. The Swin-based PIS achieves 79.01% F1, outperforming all ViT variants and underscoring hierarchical features’ importance, despite its non-temporal pretraining. AgriFM dominates with 81.90% F1 (84.63% recall), affirming both Video Swin’s architectural superiority for crop mapping and the added value of temporal-aware pretraining.
Figure 5: Comparative performance on cropland mapping and boundary delineation task. Bar plots show F1-score for the positive class across models, grouped by architecture type (CNN variants, ViT-based, and Video Swin-based). Radar plots show metrics (precision, recall, F1-score and OA) comparison across models.
Boundary delineation (Table 4) further accentuates methodological differences. ViT performance plummets (SatMAE: 62.50% F1; Prithvi: 53.45% F1), consistent with Figure 6’s oversmoothed edges—ViT’s aggressive downsampling struggles with sub-pixel boundary variations. CNN-LSTM maintains relative robustness (72.20% F1). PIS shows architectural promise (71.20% F1) but trails AgriFM by 2.58%, primarily due to lacking temporal pretraining. Intriguingly, randomly initialized Swin-scratch achieves 72.13% F1, merely 1.65% below AgriFM, suggesting hierarchical backbone inherently excels at spatial sensitivity while temporal pretraining provides systematic refinement. The underperformance of PIS versus Agri-scratch highlights incompatibilities when rearranging non-temporal pretrained weights, justifying our dedicated spatiotemporal pretraining strategy.
Figure 6: Visual comparison of cropland mapping and field boundary delineation results (green: cropland pixels, red: delineated boundaries) demonstrating method-specific spatial characteristics.
@@
Figure 7: The generated map of cropland and its boundaries in the ARA region of France for the year 2020: a) labels from EuroCrops dataset; b) the mapping results from our method. The green areas represent cropland, while the red areas delineate the boundaries of this cropland.
@@
4 Early-season Crop Mapping
This experiment specifically assesses the model’s capability to recognize crop types during early growth stages—a critical requirement for proactive agricultural management. Leveraging the same Sentinel-2 dataset (10m resolution, 10 spectral bands) and spatial coverage (ARA region, France) as the cropland mapping task, we specifically target classification of 16 major crops during their initial six-month growth phases using 2018-2020 data. The temporal sequence construction adopts a stratified sampling approach: four observations per month are selected, generating 24 temporal observations that capture early phenological development. The 2,327 tiles per year maintain strict temporal isolation: 2018 (training), 2019 (validation), and 2020 (testing).
As shown in Table 5, we systematically compare various methods on 16-crop early-season crop mapping using four metrics: precision, recall, F1-score, and overall accuracy (OA). Figure 8 presents three complementary visualizations: (a) bar charts comparing F1 scores, (b) radar plots showing metric balance, and (c) line graphs analyzing F1-scores across crops ordered by training sample frequency (low to high). Figure 9 further displays spatial mapping comparisons in representative regions.
Table 5: Performance comparison on early-season crop mapping (%).
@@ 4 Early-season Crop Mapping
The results demonstrate AgriFM’s clear superiority with 59.88% F1 and 77.51% OA. Architecturally similar approaches (PIS: 54.51% F1; AgriFM-scratch: 54.82% F1) also significantly outperform alternatives. Notably, the performance gaps between CNN/ViT methods and AgriFM (13-15% F1) are markedly larger than in previous cropland mapping tasks, underscoring the heightened challenge of early-season classification requiring precise spatiotemporal feature coordination. Three key observations emerge: (1) CNN variants (best: CNN-LSTM 44.29% F1) excel at spatial feature fusion but lack temporal modeling capacity; (2) ViT-based methods (best: SatMAE 46.10% F1) leverage global attention for temporal modeling but sacrifice spatial detail; (3) AgriFM uniquely integrates both capabilities through multi-source temporal modeling and pretraining.
@@
Figure 8 c) reveals distinct response patterns to class imbalance. ViT-based methods show advantages for rare crops (left side), benefiting from global temporal context capture. Conversely, CNNs outperform ViTs for dominant crops (right side), where local spatial features prove more discriminative. AgriFM maintains remarkably stable performance across all frequency categories, achieving high F1 improvement for rare crops versus suboptimal methods.
@@
Figure 8: Early-season crop type mapping analysis: (a) Method-wise metric comparison, (b) Balanced performance radar chart, (c) Per-class F1 scores ordered by training sample frequency.
@@
Spatial results in Figure 9 exhibit three characteristic patterns: (1) CNN variants preserve field boundaries but suffer misclassification; (2) ViT methods show superior class discrimination but blurred boundaries and small-field omissions; (3) AgriFM achieves both accurate crop identification and sharp boundary delineation. These findings collectively validate that hierarchical temporal modeling synergistically optimizes classification accuracy and spatial precision.
@@
Three principal conclusions emerge from the above analysis: First, early-season crop classification demands delicate balance between temporal and spatial feature extraction that conventional single-advantage architectures cannot provide. Second, multi-source temporal pretraining effectively mitigates class imbalance challenges. Third, AgriFM’s unified design delivers optimal performance across all scenarios, establishing a new paradigm for precision agriculture. These insights provide both theoretical and practical guidance for agricultural remote sensing model development.
Figure 9: Visual comparison of early-season crop mapping results.
@@
We also employ our trained AgriFM to generate a map of the early crop distribution in the ARA region of France for the year 2020, as depicted in Figure 10. Our model leverages satellite data from the first half of the year, ensuring an accurate mapping of crop distribution in the first half of each year. The map produced by our method effectively reflects the actual crop planting scenario for that year. As can be observed from the figure, there is a high level of crop diversity within the same area, yet our method successfully differentiates between different crops.
Figure 10: Generated map of early-season crops in the ARA region of France for the year 2020: a) labels from EuroCrops dataset; b) the mapping results from our method.
@@ 5 Paddy Rice Mapping in Monsoon Asia
This experiment evaluates the AgriFM’s capacity to synergize multi-resolution satellite data for paddy rice mapping in Asia. We fuse Harmonized Landsat-Sentinel (HLS30) imagery (30m spatial resolution, 6 spectral bands, 5 temporal steps) with MODIS observations (250m&500m resolution, 7 bands, 44 temporal steps) across Monsoon Asia’s rice cultivation belt during 2019. A total of 3,039 training samples 1013 validation samples and 1013 test samples are drawn from high-confidence rice parcels identified through consensus of existing products. Input sequences span the critical rice-growing period (January-June).
The experimental results in Table 6 demonstrate the comparative performance of different methods for paddy rice mapping across monsoon Asia using HLS30 data. All models are tested under consistent conditions, with AgriFM evaluated in two configurations: HLS30-only input and combined HLS30-MODIS input. Figure 11 shows some representative rice mapping results.
Swin-based methods consistently outperform other architectures, with PIS achieving 86.49% F1 and 90.32% average F1. This 3-4% advantage over CNN variants (best: CNN at 85.85% F1) and 4-5% over ViT models (best: SatMAE at 83.22% F1) confirms the superiority of hierarchical feature extraction for rice mapping. Notably, the randomly initialized Swin-scratch shows competitive performance (85.25% F1), suggesting the inherent suitability of Swin’s architecture for agricultural patterns.
Table 6: Comparison results on paddy rice mapping in Monsoon Asia with different methods.
Model Source Positive (Rice) Average
@@
CNN HLS30 82.61 89.35 85.85 89.04 90.82 89.85 91.43
CNN-LSTM HLS30 81.87 88.62 85.11 88.52 90.28 89.32 90.97
3DCNN HLS30 81.48 88.38 84.79 88.27 90.07 89.08 90.77
SatMAE (ViT-L) HLS30 80.72 85.88 83.22 87.39 88.73 88.01 89.92
Prithvi (ViT-b) HLS30 79.45 86.10 82.64 86.77 88.48 87.54 89.47
PIS (Swin-b) HLS30 83.46 89.75 86.49 89.56 91.22 90.32 91.84
AgriFM-scratch HLS30 80.84 90.17 85.25 88.30 90.70 89.35 90.92
AgriFM HLS30 83.30 88.98 86.04 89.32 90.82 90.02 91.60
HLS & MODIS 83.74 90.85 87.15 89.92 91.80 90.78 92.20
Figure 11: Visual comparison of paddy rice mapping in Monsoon Asia.
Considering the frequent cloud cover in the Monsoon Asia region, we incorporate MODIS data to aid in rice mapping. The experimental results suggest that the inclusion of MODIS data can enhance recall from 88.98% to 90.85%. This indicates that a larger proportion of rice can be correctly identified, emphasizing the effectiveness of our approach in extracting and integrating features from the MODIS input with those derived from HLS30. However, it can be observed that the precision decreases slightly with the inclusion of MODIS data. Although the F1 score still improves, the increase is not substantial. This is likely due to the significant difference in resolution between MODIS and HLS30. Furthermore, the accuracy of rice mapping using only HLS30 is already quite high, implying that in most scenarios, the best results can be achieved using only HLS30. Nevertheless, these findings still demonstrate that our proposed method can effectively integrate and process multi-source remote sensing data.

### v3

@@ 2.3.2 Paddy Rice Mapping in Monsoon Asia
2.3.3 Winter Wheat Mapping in Asia
@@
2.3.4 Multi-source Mapping Tasks
To evaluate the cross-spatial generalization capabilities of AgriFM, we construct a multi-source dataset incorporating Sentinel-2, Landsat 8/9, and MODIS satellite imagery with varying temporal configurations. For Sentinel-2, we utilize 32 temporal observations for agricultural land mapping and field boundary delineation, and 24 observations for agriculture land use/cover mapping. Landsat 8/9 data contained 12 frames, while MODIS provided 44 temporal observations. This configuration enables systematic analysis of model performance across different spatial resolutions and temporal densities.
The dataset design specifically addresses cross-spatial generalization by including both single-source and multi-source combinations. We evaluate pure Sentinel-2 scenarios against Landsat-only, Landsat-MODIS, Landsat-Sentinel fusion setups, allowing direct comparison of how different spatial and temporal characteristics affect mapping accuracy across all agricultural tasks. This structured approach provides comprehensive insights into the model’s ability to leverage complementary information from multi-resolution satellite systems.
2.3.5 Validation
@@
To comprehensively evaluate AgriFM’s performance, we conduct comparisons across three distinct methodological categories. First, we examine conventional deep learning architectures widely adopted in various mapping tasks: (1) A CNN (VGG16 variant (Simonyan and Zisserman, 2015)) treating temporal observations as additional input channels; (2) A CNN-LSTM hybrid extracting spatial features per timestep using VGG16 followed by temporal modeling through LSTM; and (3) A 3D CNN processing spatiotemporal cubes (height × width × time × bands). All baselines employ identical decoder structures and training protocols to isolate architectural differences, ensuring fair comparison of their spatiotemporal representation capabilities. While more advanced CNN variants (e.g., ResNet50 (He et al., 2016)) exist, VGG16 provides a balanced representation of CNN capabilities through its straightforward yet effective hierarchical feature extraction design. This choice allows us to focus on comparing fundamental architectural paradigms, CNNs versus transformer-based foundation models, rather than optimizing performance through model-specific refinements.
Second, we compare against ViT-based foundation models that incorporate temporal information during pretraining. Specifically, we evaluate SatMAE (Cong et al., 2022) (using Landsat-8 data) and Prithvi (Jakubik et al., 2023) (trained on HLS30 data), which represent the most temporally-aware open-source RSFMs currently available, despite their limited sequence lengths. In addition, we also incorporate some recently released RSFMS, Galileo (Tseng et al., 2025) and SMARTIES (Sumbul et al., 2025) to verify our method’s effeveness. These comparisons reveal how feature fusion strategies and hierarchical representations impact mapping precision, particularly for crops with subtle spectral-temporal signatures.
Third, we analyze PIS (An et al., 2024) and GFM (Mendieta et al., 2023), two Swin Transformer-based foundation models lacking explicit temporal processing. While sharing similar backbone architecture and pretraining data scales with our approach, their non-temporal designs help isolate the benefits of our temporal-aware pretraining strategy. Although they use the Swin transformer as its backbone, we can reuse some weights as pre-trained weights of the video Swin transformer and adopt the same architecture as our method for comparison.
2.3.6 Implementation details
@@
For downstream mapping tasks, we employ the same network structure, that is, the unified mapping framework that we construct. The learning rate is set to 6e-5. All experiments are conducted on four L40 GPUs, with a batch size of 16, and last for 50 epochs. We select the model that performs best on the validation dataset as our final result.
3 Agricultural Land Mapping and Field Boundary Delineation
@@
As detailed in Tables 3 and 4, we systematically evaluate three methodological categories: (1) CNN variants (CNN, CNN-LSTM, 3DCNN) representing conventional deep learning approaches; (2) ViT-based methods (SatMAE, Prithvi, Galileo and SMARTIES) as current mainstream remote sensing foundation models pretrained with temporal data; and (3) Swin transformer-based methods (PIS, GFM, AgriFM-scratch) demonstrating hierarchical feature extraction’s impact on agriculture mapping. Notably, backbones of PIS and GFM are Swin transformers and don’t adopt temporal data. We adaptively rearrange its partial parameters to initialize our Video Swin Transformer, enabling comparison of temporal-agnostic pretraining effects. AgriFM-scratch serves as the randomly initialized control. The ”Positive” column specifies agricultural land / boundary-specific metrics, while ”Average” incorporates background class performance. Figure 5 visually contrasts core metrics (Precision / Recall / F1) across methods, and Figure 7 showcases regional mapping outputs where green area denotes cropland and red area indicates boundaries.
Table 3: Performance comparison on the agricultural land mapping task (%). The positive column means metrics of agricultural land identification, while the average columns means the average metric of both background and agricultural land.
Model Positive Average
@@
AgriFM 84.31 81.90 83.09 84.58 84.42 84.48 84.61
Table 4: Performance comparison on the field boundary delineation task (%). The positive column means metrics of field boundary identification, while the average columns means the average metric of both background and field boundary.
Model Positive Average
@@
AgriFM 75.11 77.47 76.27 80.61 80.98 80.78 81.84
For agricultural land mapping (Table 3), CNN architectures demonstrate strong baseline performance, with CNN-LSTM leading at 80.39% F1-score due to its explicit spatiotemporal fusion design. 3DCNN achieves only 77.93% F1 despite 3D convolutions. This shows that simply adding 3D convolution does not guarantee better spatiotemporal modeling, and also shows that the 3D convolution in traditional deep learning methods for processing spatiotemporal data may have insufficient information extraction capabilities for mapping tasks. ViT-based foundation models trail slightly at 73.69%, 75.73%, 76.80% and 74.76% respectively, confirming ViT’s architectural constraints. The 16×16 patch embedding causes irreversible spatial detail loss (evident in Figure 7’s blurred boundaries). While increasing input resolution could theoretically mitigate this issue, the computational cost becomes prohibitive for temporal sequences while distorting original resolution relationships. The Swin-based PIS and GFM achieves 79.01% and 81.06% F1, outperforming all ViT variants and underscoring hierarchical features’ importance, despite its non-temporal pretraining. AgriFM dominates with 83.09% F1, affirming both Video Swin’s architectural superiority for agriculture mapping and the added value of temporal-aware pretraining.
Figure 5: Comparative performance on agricultural land mapping and boundary delineation task. Bar plots show F1-score for the positive class across models, grouped by architecture type (CNN variants, ViT-based, and Swin-based). Radar plots show metrics (precision, recall, F1-score and OA) comparison across models.
Boundary delineation (Table 4) further accentuates methodological differences. ViT-based methods performance plummets (Prithvi: 53.45% F1; Galileo: 62.51%; SatMAE: 62.50% F1; SMARTIES: 59.92% F1), consistent with Figure 6’s oversmoothed edges—ViT’s aggressive downsampling struggles with sub-pixel boundary variations. CNN-LSTM maintains relative robustness (72.20% F1). PIS and GFM show architectural promise (71.20% F1 and 72.38% F1) but still trail AgriFM by 5.07% and 3.89%, primarily due to lacking temporal pretraining. Intriguingly, randomly initialized AgriFM-scratch achieves 72.13% F1, merely 4.14% below AgriFM, suggesting hierarchical backbone inherently excels at spatial sensitivity while temporal pretraining provides systematic refinement. The underperformance of PIS and GFM versus Agri-scratch highlights incompatibilities when rearranging non-temporal pretrained weights, justifying our dedicated spatiotemporal pretraining strategy.
Figure 6: Visual comparison of agricultural land mapping and field boundary delineation results from test dataset in $20$. Only the best-performing methods from each task are shown to ensure clarity in comparison. Cropland pixels are shown in green and field boundaries in red. Representative regions with notable differences are highlighted with yellow bounding boxes for detailed comparison.
@@
Figure 7: The generated map of agricultural land and its boundaries in the ARA region of France for the year 2020: a) labels from EuroCrops dataset; b) the mapping results from our method. The lower panel shows the complete mapping coverage across the study area, while the upper panel provides enlarged views of selected representative regions. Green areas represent identified cropland, with red boundaries delineating individual field parcels.
@@
4 Agricultural Land Use / Land Cover Mapping
This experiment specifically assesses AgriFM’s capability to recognize agricultural land use / land cover during early growth stages—a critical requirement for proactive agricultural management. Leveraging the same Sentinel-2 dataset (10m resolution, 10 spectral bands) and spatial coverage (ARA region, France) as the agricultural land mapping task, we specifically target classification of 16 major classes during their initial six-month growth phases using 2018-2020 data. The temporal sequence construction adopts a stratified sampling approach: four observations per month are selected, generating 24 temporal observations that capture early phenological development. The 2,327 tiles per year maintain strict temporal isolation: 2018 (training), 2019 (validation), and 2020 (testing).
As shown in Table 5, we systematically compare various methods with four metrics: precision, recall, F1-score, and overall accuracy (OA). Figure 8 presents three complementary visualizations: (a) bar charts comparing F1 scores, (b) radar plots showing metric balance, and (c) line graphs analyzing F1-scores across land cover / land use ordered by training sample frequency (low to high). Figure 9 further displays spatial mapping comparisons in representative regions.
Table 5: Performance comparison on agricultural land use / land cover mapping (%). Metrics are computed as average of each class.
@@ 4 Agricultural Land Use / Land Cover Mapping
The results demonstrate AgriFM’s clear superiority with 60.49% F1 and 77.38% OA. Architecturally similar approaches (PIS: 54.51% F1; GFM: 57.75%; AgriFM-scratch: 54.82% F1) also significantly outperform alternatives. Notably, the performance gaps between CNN / ViT methods and AgriFM (13-15% F1) are markedly larger than in the agricultural land mapping task, underscoring the heightened challenge of agricultural land use / land cover mapping requiring precise spatiotemporal feature coordination. Three key observations emerge: (1) CNN variants (best: CNN-LSTM 44.29% F1) excel at spatial feature fusion but lack temporal modeling capacity; (2) ViT-based methods (best: SMARTIES 47.65% F1) leverage global attention for temporal modeling but sacrifice spatial detail; (3) AgriFM uniquely integrates both capabilities through multi-source temporal modeling and pretraining.
@@
Figure 8 c) reveals distinct response patterns to class imbalance. ViT-based methods show advantages for rare classes (left side), benefiting from global temporal context capture. Conversely, CNNs outperform ViTs for dominant classes (right side), where local spatial features prove more discriminative. AgriFM maintains remarkably stable performance across all frequency categories, achieving high F1 improvement for rare classes versus suboptimal methods.
@@
Figure 8: Agricultural land use / land cover mapping analysis: (a) Method-wise metric comparison, (b) Balanced performance radar chart, (c) Per-class F1 scores ordered by training sample frequency.
@@
Spatial results in Figure 9 exhibit three characteristic patterns: (1) CNN variants preserve field boundaries but suffer misclassification; (2) ViT methods show superior class discrimination but blurred boundaries and small-field omissions; (3) AgriFM achieves both accurate identification and sharp boundary delineation. These findings collectively validate that hierarchical temporal modeling synergistically optimizes classification accuracy and spatial precision.
@@
Three principal conclusions emerge from the above analysis: First, agricultural land use / land cover demands delicate balance between temporal and spatial feature extraction that conventional single-advantage architectures cannot provide. Second, multi-source temporal pretraining effectively mitigates class imbalance challenges. Third, AgriFM’s unified design delivers optimal performance across all scenarios, establishing a new paradigm for precision agriculture. These insights provide both theoretical and practical guidance for agricultural remote sensing model development.
Figure 9: Visual comparison of agricultural land use / land cover mapping results from test dataset in $20$. Only the best-performing methods from each task are shown to ensure clarity in comparison. Different colors in the bottom legend reprensent different land use / land covers. Representative regions with notable differences are highlighted with read bounding boxes for detailed comparison.
@@
We also employ our trained AgriFM to generate a map of agricultural land use / land cover in the ARA region of France for the year 2020, as depicted in Figure 10. Our model leverages satellite data from the first half of the year, ensuring an accurate mapping of crop distribution in the first half of each year. The map produced by our method effectively reflects the actual crop planting scenario for that year. As can be observed from the figure, there is a high level of crop diversity within the same area, yet our method successfully differentiates between different land use / land covers.
Figure 10: Generated map of agricultural land use / land cover mapping in the ARA region of France for the year 2020: a) labels from EuroCrops dataset; b) the mapping results from our method. The lower panel shows the complete mapping coverage across the study area, while the upper panel provides enlarged views of selected representative regions.
@@ 5 Paddy Rice Mapping in Monsoon Asia
This experiment evaluates AgriFM’s capacity to utilize temporal satellite data for paddy rice mapping in Asia. We utilize Harmonized Landsat-Sentinel (HLS30) imagery (30m spatial resolution, 6 spectral bands, 5 temporal steps) across Monsoon Asia’s rice cultivation belt during 2019. A total of 3,039 training samples, 1,013 validation samples, and 1,013 test samples are drawn from high-confidence rice parcels identified through consensus of existing products. Input sequences span the critical rice-growing period (January–June).
The experimental results in Table 6 demonstrate the comparative performance of different methods for paddy rice mapping across Monsoon Asia using HLS30 data. All models are tested under consistent conditions. Figure 11 shows some representative rice mapping results.
Swin-based methods consistently outperform other architectures, with PIS achieving 86.49% F1 and GFM 86.31% F1. This advantage over CNN variants (best: CNN at 85.85% F1) and ViT models (best: Galileo at 84.79% F1) confirms the superiority of hierarchical feature extraction for rice mapping. Notably, the randomly initialized Swin-scratch shows competitive performance (85.25% F1), suggesting the inherent suitability of Swin’s architecture for agricultural patterns.
Table 6: Performance comparison on the paddy rice mapping task (%). The positive column means metrics of paddy rice identification, while the average columns means the average metric of both background and paddy rice.
Model Positive (Paddy Rice) Average
@@
CNN 82.61 89.35 85.85 89.04 90.82 89.85 91.43
CNN-LSTM 81.87 88.62 85.11 88.52 90.28 89.32 90.97
3DCNN 81.48 88.38 84.79 88.27 90.07 89.08 90.77
Prithvi (ViT-b) 79.45 86.10 82.64 86.77 88.48 87.54 89.47
Galileo (ViT-b) 82.12 87.63 84.79 88.68 90.23 89.94 91.42
SatMAE (ViT-L) 80.72 85.88 83.22 87.39 88.73 88.01 89.92
SMARTIES (ViT-L) 81.37 81.01 84.10 88.18 89.76 88.92 91.02
PIS (Swin-b) 83.46 89.75 86.49 89.56 91.22 90.32 91.84
GFM (Swin-b) 83.12 89.75 86.31 89.38 91.13 90.18 91.71
AgriFM-scratch 80.84 90.17 85.25 88.30 90.70 89.35 90.92
AgriFM 84.03 90.12 86.97 89.93 91.54 90.67 92.14
Further analysis reveals that AgriFM achieves the highest overall performance, with an F1 score of 86.97% for paddy rice mapping, outperforming all other methods including state-of-the-art Swin-based models like PIS and GFM. This improvement can be attributed to AgriFM’s ability to effectively leverage multi-source temporal data and its synchronized spatio-temporal downsampling strategy, which enhances feature representation for complex phenological patterns. Specifically, AgriFM excels in precision (84.03%) while maintaining high recall (90.12%), indicating a better balance in minimizing false positives and capturing true rice parcels. The superior performance across all average metrics (precision: 89.93%, recall: 91.54%, F1: 90.67%, OA: 92.14%) underscores its robustness and generalizability in diverse rice-growing regions. Compared to AgriFM-scratch, which relies solely on architectural advantages, the pre-trained AgriFM demonstrates the added value of land cover fraction supervision and pre-training, leading to more reliable mappings. These results highlight AgriFM’s potential as a versatile foundation model for precise agricultural monitoring in challenging, large-scale environments.
Figure 11: Visual comparison of paddy rice mapping results in Monsoon Asia from test dataset in 2019. Only the best-performing methods from each task are shown to ensure clarity in comparison. Paddy rice pixels are shown in green. Representative regions with notable differences are highlighted with red bounding boxes for detailed comparison.